# SP2T: Sparse Proxy Attention for Dual-stream Point Transformer

### v1

@@ 3.3 SPA for Point-Proxy Interaction
Figure 5: Overview of SPA calculation. Map-reduce computing based on association tuple lists are utilized by SPA to compute sparse attention between proxy and point.
@@
Sparse Proxy Attention (SPA). Unlike traditional attention [38], the attention between the point and the proxy is sparse and discrete due to the vertex-based connections, leading to substantial computational demands with the typical attention operator. To address this issue, we implemented sparse proxy attention to speed up information exchange between proxies and points. In practice, sparse proxy attention utilizes either the point cloud or proxy features as the query feature, while the other acts as the key and value features. Thus, it enables the aggregation of information from the point to the proxy and the feedback from the proxy back to the point.
@@
Take the SPA that transfers features from proxy to point as an example. As shown in Fig. 5, to calculate the similarity between the proxy and the point, we denote the indices of the corresponding original points and proxy points in the $i$-th association tuple as $a^{pt}_{i},a^{px}_{i}$. The sparse attention module first computes the exponential similarity for the query of each association group and the key features of each head.
$S^{h}_{i}=\exp\left(\mathop{dot}\left(q^{h}_{a^{pt}_{i}},k^{h}_{a^{px}_{i}}\right)/\sqrt{d}\right)$ (1)
where $S^{h}_{i}$ is the similarity of the $i$-th association of the head $h$, $q^{h}_{a^{pt}_{i}}$ and $k^{h}_{a^{px}_{i}}$ is the query $h$ of the head and key of $a^{pt}_{i},a^{px}_{i}$. $d$ is the dimension of the feature.
@@
Building on this, we then employ a map-reduce algorithm to obtain the sum of exponentiated similarity over each query. Finally, a division is performed to complete the sparse softmax and obtain the weights for each value.
$W^{h}_{i}=\frac{S^{h}_{i}}{\sum_{j,a^{px}_{j}=a^{px}_{i}}S^{h}_{j}}$ (2)
@@
where $S^{h}_{i},S^{h}_{j}$ is the similarity of the $i$-th association and the $j$-th association of the same proxy. $W^{h}_{i}$ is the weight of the value after the sparse softmax.
Figure 6: Overview of Table-based Relative Bias (TRB). TRB improves the spatial understanding of subsequent attention by sampling the Relative Bias Table, thus obtaining the relative bias linked to that relative distance.
@@
Finally, we adopt a similar map-reduce-based method to weight the corresponding value features by the weights and association indices to obtain the output features.
$o^{h}_{i}=\sum_{j,a^{px}_{j}=i}v^{h}_{a^{pt}_{j}}W^{h}_{i}$ (3)
@@
where $v^{h}_{a^{pt}_{j}}$ is the association value $a^{pt}_{j}$ of the $h$ th head and $o^{h}_{i}$ is the attention output of the $h$ th head proxy $i$.
@@
Table-based Relative Bias. Inspire by the Swin Transformer [24], we found that integrating relative bias (RB) into the attention mechanism improves the network’s proficiency in capturing the connections between the proxy and the point cloud. However, the relative bias in Swin Transformer is a learning parameter synchronized with the attention map, while in SPA, the relative bias must adapt to match the dimensions of the point cloud and proxy.
@@
To address the aforementioned issues, we introduce Table-Based Relative Bias (TRB). As shown in Fig. 6, the TRB utilizes a lookup table $T_{rpe}$ characterized by dimensions $X_{rpe},Y_{rpe},Z_{rpe}$. Initially, the input relative positions are converted into normalized coordinates for the lookup table by applying a scaling factor $s_{rpe}$. The biases are then derived at these coordinates via trilinear interpolation. Relative distances that are beyond the permissible range are clamped to valid numerical values. In terms of form, the formula for TRB is as follows.
$TRB(x)=\operatorname{TGS}\left(T_{rpe},\mathop{clamp}(s_{rpe}x,-1,1)\right)$ (4)
@@
Finally, the similarity formula with TRB is shown as follows.
$S^{h}_{i}=\exp\left(\mathop{dot}\left(q^{h}_{a^{pt}_{i}},k^{h}_{a^{px}_{i}}\right)/\sqrt{d}\right)+\operatorname{TRB}^{h}\left(p^{pt}_{a^{pt}_{i}}-p^{px}_{a^{px}_{i}}\right)$ (5)
where $S^{h}_{i}$ is the similarity of the $i$-th association of the head $h$, $q^{h}_{a^{pt}_{i}}$ and $k^{h}_{a^{px}_{i}}$ is the query of the head $h$ and key of $a^{pt}_{i},a^{px}_{i}$. $d$ is the dimension of the feature. $p^{pt}_{a^{pt}_{i}},p^{px}_{a^{px}_{i}}$ is the position of the point and proxy.
Furthermore, the feature dimension of the lookup table $T_{rpe}$ is configured to match the number of heads in the sparse attention mechanism. This setup allows parallel execution of interpolation sampling and integration into similarity for each head. Furthermore, during training, the initialization of $T_{rpe}$ used a Gaussian distribution with scaled variance, which encourages the network to initially focus on the features that are proximate.
During the implementation phase, the point and proxy positions remain unchanged within the same layer. This means that for each instance of sparse attention in this layer, the relative position input provided to the relative position encoding module remains constant. Consequently, it is possible to compute the relative bias for a layer with a single invocation. Building on this optimization, we share all relative bias values across each layer, thus reducing model complexity and computational demand. Additionally, since the proxy positions are constant throughout the network, we also test sharing the relative bias amongst proxies across the entire network and across different layers.
@@
SP2T 78.7 74.9 37.0 35.2 73.4 77.9
@@
Method Metric Area1 Area2 Area3 Area4 Area5 Area6 6-Fold
PTv2 [43] allAcc 92.30 86.00 92.98 89.23 91.24 94.26 90.76
mAcc 88.44 72.81 88.41 82.50 77.85 92.44 83.13
mIoU 81.14 61.25 81.65 69.06 72.02 85.95 75.17
PTv3 [44] allAcc 93.22 86.26 94.56 90.72 91.67 94.98 91.53
mAcc 89.92 74.44 94.45 81.11 78.92 93.55 85.31
mIoU 83.01 63.42 86.66 71.34 73.43 87.31 77.70
SP2T allAcc 92.69 87.96 93.43 90.31 91.87 94.81 91.85
mAcc 90.21 77.74 93.67 81.85 79.24 93.48 86.03
mIoU 83.27 64.87 86.12 71.89 73.43 87.84 77.90
Table 2: S3DIS 6-fold cross-validation.
3.4 Local and Global Fusion
Local Fusion The local fusion in SP2T is followed by PTv3 [44]. In particular, SP2T employs point-cloud serialization along with self-attention within PTV3 to facilitate the local fusion of point-cloud data.
Global Fusion Fig. 7 (b) illustrates the global fusion using a self-attention module with table-based relative bias. The proxy interaction module facilitates the exchange of information across all proxies, allowing each to have direct access to the global context.
4 Experiments
4.1 Result of Downstream Tasks
In this section, we will compare our model with other SOTA models in various downstream tasks. For detailed model and training settings, we report them in supplementary material.
Indoor Semantic Segmentation. In Tab. 2, we present the validation and test results with other methods [43, 44, 50, 40, 31, 7] on the ScanNet v2 [10] and ScanNet200 [32] benchmarks, as well as Area 5 and the 6-fold cross-validation [29] conducted on S3DIS [1] (see Tab. 2 for detailed information). Our model demonstrated a 1.3% / 1.3% improvement in the ScanNet Val / Test set and showed an 1.8% / 1.2% increase in performance in the ScanNet200 Val / Test set compared to PTv3 [44]. In S3DIS, our performance remained consistent with PTv3 in S3DIS area5, and has a 0.2% increase in performance in S3DIS 6-fold.
As noted in the PTv3 issues on GitHub, compared to the configuration in the val-set, the PTv3 test results could have used extra test time augmentation (TTA), such as incorporating data from the validation set for training, combining results from multiple training models [20], and applying over-segmentation methods [27, 33, 45]. These augmentations may not be open source in the code of PTv3. To ensure a fair comparison, our model compares the PTv3 test results with the TTA of the val set. Furthermore, we provide more details of the test result in the supplementary material.
@@ 4.1 Result of Downstream Tasks
Indoor instance Segmentation. In Tab. 3, we provide validation outcomes for our model on the instance segmentation benchmarks of ScanNet v2 [10] and ScanNet200 [32]. Performance metrics displayed include mAP, mAP25, and mAP50, and are compared to various state-of-the-art (SOTA) backbones. Following PTv3 [44], we use the instance segmentation framework by utilizing PointGroup [19] in all evaluations, altering only the backbone. Compared to PTv3, we achieved enhancements of 1.1% in mAP25, 0.6% in mAP50, and 0.2% in mAP on ScanNet. Furthermore, in ScanNet200, there were gains of 1.1% in mAP25, 0.9% in mAP50, and 0.2% in mAP.
@@ 4.2 Ablation Study
Followed by PTv3, the model’s ablation studies were conducted on the ScanNet validation set, where four metrics, including mIOU, mAcc, allAcc and latency. The latency is reported using a single 4090 GPU with 1 batch size.
The number of proxy and association. Tab. 4 presents the ablation study exploring the impact of the number of proxies and associations in SP2T. Experimental results indicate that increasing the number of proxies and associations does not always result in better performance. On the one hand, an excess of proxies leads to a shortage of corresponding point clouds, thereby restricting the model’s receptive field. On the other hand, an abundance of associations generates numerous point clouds linked to the proxies, making it difficult for the proxy to accurately represent the point cloud. As a result, the ideal number of proxies and associations is essential to produce the best results.
Number Method Metrics
Proxy Association Association mIou mAcc allAcc Time/ms
Baseline (PTV3) 77.51 85.16 92.24 61
40 8 Vertex 77.74 86.39 92.05 65
80 8 Vertex 77.79 86.11 92.17 68
120 8 Vertex 77.71 86.22 92.43 71
160 8 Vertex 78.71 86.23 92.51 74
320 8 Vertex 78.08 86.15 92.31 92
512 8 Vertex 78.22 86.60 92.30 120
800 8 Vertex 77.79 85.83 92.01 221
160 4 KNN 77.99 86.61 92.22 69
160 8 KNN 78.71 86.23 92.51 83
160 12 KNN 77.90 85.77 92.19 97
160 16 KNN 78.10 86.99 92.40 158
Table 4: Ablation study about the number of proxy and association of SP2T in ScanNet. KNN utilizes the L-Infinity norm for measuring distance.
TRB FPS Fix Number Fix Size Spatial-wise
@@
Table 5: Ablation study about proxy sample method and TRB in ScanNet. OOM means Out of Memory.
Share TRB mIoU mAcc allAcc Time
@@
Table 6: Ablation study about parameter sharing of TRB.
Proxy Sampling Method and Table-based Relative Bias. Tab. 5 provides the results of the ablation study on proxy sampling methods and Table-based Relative Bias. The experiment indicates that the spatial-wise sampling method yields the best model accuracy among various sampling methods. Additionally, TRB is adaptable to all sampling methods and can generally enhance model performance. Particularly with respect to the sampling method, the Fix Number and Fix Size-based methods struggle with point clouds at different scales, resulting in reduced accuracy and failure training. In contrast, the spatial-wise sampling approach can flexibly determine the proxy spacing thereby improving performance. Regarding the FPS-based method, the analysis is provided in the Discussion regarding the low performance of the FPS-based sampling method.
Share of Table Relative Bias. Tab. 6 shows a comparison of the accuracy of the model w/ and w/o the shared TRB. The experiment shows that the shared TRB both improves the accuracy of the model and reduces the inference time. Furthermore, Fig. 8 shows the MSE distance for TRB during point-proxy and proxy-proxy interaction in different layers. In total, TRB demonstrates stage-level similarity in the proxy-proxy interaction, while all TRB is similar in the point-proxy interaction. Consequently, the sharing of TRB improves model accuracy and reduces inference time.
Empty Proxies. Different from image, grid sampling in point cloud unavoidably results in empty proxies, which lack any point cloud association. Tab. 7 presents a comparison of SP2T accuracy both with and without considering empty proxies during global fusion. The experimental data indicate that empty proxies actually enhance the accuracy of the model. We believe that the experiment proves that ‘nothing’ is also a very important piece of information in the point cloud. Although empty proxies lack point-cloud data, the absence itself conveys spatial structure information, thereby contributing to an understanding of overall spatial feature for point cloud.
Figure 8: (a) MSE of TRB in different proxy-point interaction layer. (b) MSE of TRB in different proxy-proxy interaction layer.
Empty Proxy mIoU mAcc allAcc

### v2

@@ 3.3 SPA for Point-Proxy Interaction
Sparse Proxy Attention (SPA). Unlike traditional attention [50], the attention between the point and the proxy is sparse and discrete due to the vertex-based connections, leading to substantial computational demands with the typical attention operator. We implemented sparse proxy attention to address this issue and speed up information exchange between proxies and points. In practice, sparse proxy attention utilizes the points or proxy features as the query feature, while the other acts as the key and value features. Thus, it enables the aggregation of information from the point to the proxy and the feedback from the proxy back to the point.
@@
Take the SPA that transfers features from proxy to point as an example. As shown in Fig. 5, to calculate the similarity between the proxy and the point, we denote the indices of the corresponding original points and proxy points in the $i$-th association tuple as $as^{pt}_{i},as^{px}_{i}$. The sparse attention module first computes the exponential similarity for each association group’s query and each head’s key features.
$S^{h}_{i}=\exp\left(q^{h}_{as^{pt}_{i}}\cdot k^{h}_{as^{px}_{i}}/\sqrt{d}\right)$ (1)
where $S^{h}_{i}$ is the similarity of the $i$-th association of the head $h$, $q^{h}_{as^{pt}_{i}}$ and $k^{h}_{as^{px}_{i}}$ is the query $h$ of the head and key of $as^{pt}_{i},as^{px}_{i}$. $d$ is the dimension of the feature.
@@
Building on this, we employ a map-reduce algorithm to obtain the sum of exponential similarity over each query. Finally, a division is performed to complete the sparse softmax and get the weights for each value.
Figure 5: Overview of SPA calculation. SPA utilizes map-reduce computing based on association tuple lists to compute sparse attention between proxy and point.
$W^{h}_{i}=\frac{S^{h}_{i}}{\sum_{j,as^{px}_{j}=as^{px}_{i}}S^{h}_{j}}$ (2)
@@
where $S^{h}_{i},S^{h}_{j}$ is the similarity of the $i$-th association and the $j$-th association of the same proxy. $W^{h}_{i}$ is the weight of the value after the sparse softmax. $as^{px}_{i},as^{px}_{j}$ represents the index of the association proxy for points $i$ and $j$.
@@
Finally, we adopt a similar map-reduce-based method to weigh the corresponding value features by the weights and association indices to obtain the output features.
$o^{h}_{i}=\sum_{j,as^{px}_{j}=i}v^{h}_{as^{pt}_{j}}W^{h}_{i}$ (3)
@@
where $v^{h}_{as^{pt}_{j}}$ is the association value $as^{pt}_{j}$ of the $h$-th head and $o^{h}_{i}$ is the attention output of the $h$-th head proxy $i$. The computational complexity of SPA is $O(knd)$, where $k$ is the association number and equal to 8, $n$ is the number of points, and $d$ is the channel dimension.
@@
Table-based Relative Bias. Inspired by the Swin Transformer [31], we found that integrating relative bias (RB) into the attention mechanism improves the network’s proficiency in capturing the connections between the proxy and the points. However, the relative bias in the Swin Transformer is a learning parameter synchronized with the attention map. In contrast, in SPA, the relative bias must adapt to match the dimensions of the points and proxy.
@@
We introduce Table-Based Relative Bias (TRB) to address the abovementioned issues. The TRB utilizes a lookup table $T_{rpe}$ characterized by dimensions $X_{rpe},Y_{rpe},Z_{rpe}$. The input relative positions are initially converted into normalized coordinates for the lookup table by applying a scaling factor $s_{rpe}$. The biases are then derived at these coordinates via trilinear interpolation. Relative distances beyond the permissible range are clamped to valid numerical values. In terms of form, the formula for TRB is as follows.
$\operatorname{TRB}(x)=\operatorname{TGS}\left(T_{rpe},\operatorname{clamp}(s_{rpe}x,-1,1)\right)$ (4)
@@
Finally, the similarity formula with TRB is shown below.
$S^{h}_{i}=\exp\left(q^{h}_{as^{pt}_{i}}\cdot k^{h}_{as^{px}_{i}}/\sqrt{d}\right)+\operatorname{TRB}^{h}\left(p^{pt}_{as^{pt}_{i}}-p^{px}_{as^{px}_{i}}\right)$ (5)
where $S^{h}_{i}$ is the similarity of the $i$-th association of the head $h$, $q^{h}_{as^{pt}_{i}}$ and $k^{h}_{as^{px}_{i}}$ is the query of the head $h$ and key of $as^{pt}_{i},as^{px}_{i}$. $d$ is the dimension of the feature. $p^{pt}_{as^{pt}_{i}},p^{px}_{as^{px}_{i}}$ is the position of the point and proxy.
Furthermore, the feature dimension of the lookup table $T_{rpe}$ is configured to match the number of heads in the sparse attention mechanism. This setup allows parallel execution of interpolation sampling and integration into similarity for each head. Furthermore, the initialization of $T_{rpe}$ used a Gaussian distribution with scaled variance, which initially encourages the network to focus on the proximate features.
@@
Outdoor Sem. Seg. nuScenes [4] Sem.KITTI [2] Waymo Val [48]
Methods Val Test Val Test mIOU mAcc
PointASNL [57] - - 48.8 - - -
MinkUNet [9] 73.3 - 63.8 - 65.9 76.6
SphereFormer [25] 78.4 81.9 67.8 74.8 69.9 -
PTv2 [55] 80.2 82.6 70.3 72.6 70.6 80.2
PTv3 [56] 80.4 82.7 70.8 74.2 71.3 80.5
SP2T 81.2 83.4 71.7 75.4 71.9 82.5
Table 1: Outdoor semantic segmentation.
Waymo [48] Det. Vehicle L2 Pedestrian L2 Cyclist L2 Mean L2
Methods # mAP mAPH mAP mAPH mAP mAPH mAPH
CenterPoint [60] 1 66.7 66.2 68.3 62.6 68.7 67.6 65.5
PillarNet [26] 1 70.4 66.9 71.6 64.9 67.8 66.7 67.2
PTv3 [56] 1 71.2 70.8 76.3 70.4 71.5 70.4 70.5
SP2T 1 72.1 71.6 76.7 71.5 74.2 73.3 72.1
CenterPoint [60] 2 67.7 67.2 71.0 67.5 71.5 70.5 68.4
PillarNet [26] 2 71.6 71.6 74.5 71.4 68.3 67.5 70.2
PTv3 [56] 2 72.5 72.1 77.6 74.5 71.0 70.1 72.2
SP2T 2 72.9 72.6 78.1 75.2 75.1 74.2 74.0
Table 2: Waymo object detection.
3.4 Local and Global Fusion
Local Fusion. The local fusion module, implemented using cross-architecture implementations (MinkUnet/Ptv3), performs local feature aggregation via point interactions. PTv3 was selected for its advantages in local fusion. In particular, SP2T employs point-cloud serialization attention to facilitate the local fusion of points.
Global Fusion. Global fusion extends proxy receptive fields through TRB-driven self-attention, with experimental validation that confirms its critical role in SP2T. Fig. 6 (b) illustrates the global fusion using a self-attention with TRB, which facilitates the exchange of information across all proxies, allowing each to have direct access to the global context.
4 Experiments
4.1 Result of Downstream Tasks
This section will compare our model with other SOTA models in various downstream tasks. For detailed model and training settings, we report them in supplementary material.
Outdoor Semantic Segmentation. The validation and test results of our model for the nuScenes [4], Semantic KITTI [2] and Waymo [48] benchmarks with other methods [9, 8, 58, 25, 55, 56, 57] are enumerated in Tab. 2. Our model outperformed existing models, obtaining the highest results for both datasets. Specifically, relative to SphereFormer, we achieved a 2.8%, 3.9%, and 2.0% boost on scenes, SemanticKITTI, and Waymo. In contrast to the proxy-based method, such as PointASNL [57], SP2T exhibits a significant improvement in SemanticKITTI by 22.9%, underscoring the critical role of sparse proxy attention.
@@ 4.1 Result of Downstream Tasks
Proxy-based Method:
PointASNL [57] 63.0 - - - 68.7 -
Fast PT [37] 72.1 - - - 70.1 -
SPoTr [38] - - - - 70.8 -
SP2T 78.7 74.9 37.0 35.2 74.6 79.7
@@
Outdoor Object Detection. In Tab. 2, we present a comparison of SP2T with other single-stage 3D detectors [60, 26, 56] on the Waymo Object Detection benchmark [48]. Our baseline is CenterPoint [60] followed by PTv3 [56]. Our model consistently surpasses sparse convolutional [60, 26] and transformer-based detectors [56], demonstrating enhancements over recent SOTA. Notably, SP2T surpasses PTv3 [56] by 1.6% using a single input frame and retains a 1.8% lead in multi-frame scenarios.
Indoor Semantic Segmentation. In Tab. 4, we compare our validation and test outcomes with other approaches [55, 56, 68, 52, 42, 9, 59] on the ScanNet v2 [11] and ScanNet200 [43] benchmarks. This includes Area 5 and the 6-fold cross-validation [40] evaluated on S3DIS [1]. SP2T achieved a performance boost of 1.2% in the ScanNet and showed enhancements of 1.8% in the ScanNet200, compared to PTv3 [56]. In S3DIS, SP2T achieved performance gains of 1.2% and 2.0% for Area 5 and 6-folds over PTv3. Compared to proxy-based approaches like PointASNL [57], Fast PT [37] and SPoTr [38], demonstrates a 15.7% over PointASNL and 5.6% over Fast PT on ScanNet, and shows gains of 5.9% over PointASNL, 4.5% over Fast PT and 3.8% over SPoTr on S3DIS, proofing the breakthrough of SP2T in proxy-based model design. In the ScanNet and ScanNet 200 test data sets, the results of the models are presented using the augmentation of test time within the val set and do not use additional methods such as over-segmentation [36, 44, 59], aggregating results from several trained models [24], or employing more data during training. More details are provided in supplementary material.
No. Methods mIOU mAcc allAcc Times/ms
1 PTv3 (Baseline) 77.51 85.16 92.24 61
2 PTv3 with 1.1x channels 77.64 ($\uparrow$ 0.13) 85.41 92.28 73
3 #1 + Pool-based PPI 77.47 ($\downarrow$ 0.04) 85.09 92.15 64
4 + Global Fusion 77.61 ($\uparrow$ 0.10) 85.08 92.07 69
5 #1 + Conv-based PPI 77.70 ($\uparrow$ 0.19) 85.87 92.20 65
6 + Global Fusion 78.17 ($\uparrow$ 0.66) 86.13 92.16 70
7 #1 + Attn-based PPI w/o TRB 77.83 ($\uparrow$ 0.32) 85.93 92.14 68
8 + Global Fusion 78.24 ($\uparrow$ 0.73) 86.22 92.27 73
9 #1 + Attn-based PPI w/ TRB 78.32 ($\uparrow$ 0.81) 86.16 92.37 69
10 + Global Fusion 78.71 ($\uparrow$ 1.20) 86.23 92.51 74
Table 5: Ablation study about structure of SP2T. PPI means Point-Proxy Interaction.
Indoor instance Segmentation. In Tab. 4, we provide validation outcomes for our model on the instance segmentation benchmarks of ScanNet v2 [11] and ScanNet200 [43]. Performance metrics displayed include mAP, mAP25, and mAP50, and are compared to various state-of-the-art (SOTA) backbones. Following PTv3 [56], we use the instance segmentation framework utilizing PointGroup [22] in all evaluations, altering only the backbone. Compared to PTv2 [55], we achieved enhancements of 2.5% in mAP25, 2.9% in mAP50, and 3.9% in mAP on ScanNet. Furthermore, in ScanNet200, there were gains of 2.0% in mAP25, 2.4% in mAP50, and 3.3% in mAP.
@@ 4.2 Ablation Study
The structure of SP2T. In Tab. 5, the structure ablation experiments of SP2T are presented. The first two rows of Tab. 5 detail the baseline PTv3 [56], along with a version of PTv3 that matches the final size of the SP2T model. This analysis shows that simply enlarging PTv3’s parameters does not significantly enhance metrics because of limited receptive fields. Therefore, SP2T’s improved performance is not due to having more parameters. The 3 and 4 rows of Tab. 5 focus on a proxy model employing pool-based proxy-point interactions (PPI), revealing that poor-based PPI between agent and point cloud can degrade performance. Furthermore, rows 5 and 6 illustrate a model with convolutional PPI via TRB, compared to one using only TRB, showing that convolution improves feature extraction in the model, unlike pooling. Finally, rows 7 to 10 of Tab. 5 examine models with and without positional coding in attentional PPI. These findings indicate that although using attention or TRB alone is beneficial, their integration notably boosts performance, aligning with previous attention studies [16, 23]. Moreover, ablation tests involving global fusion show enhanced performance across interaction modes, thanks to the global fusion capability to broadly extend the agent model’s receptive field.
The number of proxy and association. Tab. 6 rows 1 to 6 display how varying the number of proxies affects SP2T. The results indicate that the model demonstrates limited sensitivity to changes in the number of proxies, with the accuracy showing only minor fluctuations once the threshold (120) is reached. In contrast, rows 7 to 10 of Tab. 6 address the study of association counts in SP2T. The results reveal that the model reacts more notably to the different numbers of associations. Neither excessively high nor low association counts have a more pronounced impact on segmentation accuracy than variations in proxy numbers.
No. Proxy Asso. Sample E.P. mIou mAcc allAcc Time/ms